OpenStudy (anonymous):

X^2 - 4x -2 = 0 Solve by completing the square

OpenStudy (anonymous):

\[x^2-4x=2\] \[(x-2)^2=2+4=6\] \[x-2=\pm\sqrt{6}\] \[x=2\pm\sqrt{6}\]

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

yw